OpenStudy (anonymous):

Let sin x=.77 and sin y=.33 Which of the following is cos 2x +cos2y? A).62 B).60 C)-.20 D)-24

Directrix (directrix):

@MALTMAN0004 I think we will need the double angle formula for cosine. Of the three versions, go with the one marked in the attachment.

Directrix (directrix):

Let sin x=.77 and sin y=.33 cos 2x +cos2y = 1 - 2 sin²x + 1 - 2 sin²y Substitute in for sin²x and sin²y and simplify. Post what you get. @MALTMAN0004

OpenStudy (anonymous):

1.54 1-3.08+1-.66 -2.08+.34

Directrix (directrix):

I am not sure what your work represents. Or, maybe I did the wrong thing. Here's how I started the calculation. See what you think. 1 - 2(.77)² + 1 - 2(.33)² = ?

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

i'm not doing it right my answer doesn't match any of the possible answers

Directrix (directrix):

You have to crank this out: 1 - 2(.77)² + 1 - 2(.33)² = ? 1 - 2(.77)² + 1 - 2(.33)² = 1 + 1 - 2(.77)² - 2(.33)² = 2 - 2(.77)² - 2(.33)² = ? Get your calculator out and do the above operations.

Directrix (directrix):

What are the possible answers? You can always delete them after I view them.

OpenStudy (anonymous):

2-3.08-1.32

Directrix (directrix):

What does 2-3.08-1.32 equal?

OpenStudy (anonymous):

-2.4

Directrix (directrix):

That is not what this: 1 - 2(.77)² + 1 - 2(.33)² = 1 + 1 - 2(.77)² - 2(.33)² is. Is there an answer option that begins .59..?

OpenStudy (anonymous):

close.60

Directrix (directrix):

Well, look at this and see what you think: http://tinyurl.com/mv3r95f I think your calculator is messing up.

OpenStudy (anonymous):

i can't get DEG off

Directrix (directrix):

You don't need degrees. Here's how it went: Let sin x=.77 and sin y=.33 cos 2x +cos2y = 1 - 2 sin²x + 1 - 2 sin²y Substitute in for sin²x and sin²y and simplify. cos 2x +cos2y = 1 - 2(.77)² + 1 - 2(.33)² = 1 + 1 - 2(.77)² - 2(.33)² = .5964 -->

OpenStudy (anonymous):

thank you very much for the link

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

my calculator won't let me take off DEG setting

Directrix (directrix):

I don't know what kind of calculator you have but you should be able to change the DEG setting to RAD (for radians) or GRAD. DEG is the default setting on my calculator. Lately, I've been using the online calculators.

OpenStudy (anonymous):

thank you againg